# long-var-wrap = dv

### v1

@@ IV.2 Continuous Symmetry
We can find an exact solution for the $\mathrm{SO}(2)$ model in the case of $\bar{\varphi}=0$, corresponding to a slice of the phase diagram of the last section where the entire plane, for $r<0$, is unstable [54, 51]. This solution has the form
@@
Decreasing further below $r=0$ yields shorter wavelength solutions. Minimizing the free energy density $\mathcal{F}=-\frac{3}{4}uA^{2}$ corresponds to maximizing the amplitude of the wave. The global minimum is the state with the longest wavelength $k=2\pi/L$ where $L$ is the system size, while solutions with higher wave numbers are metastable states, at least for $\alpha=0$. Assuming $|r|\gg k^{2}$, the amplitude of the wave is $A\approx\sqrt{-r/u}$, corresponding to mapping the circle in field space around the bottom of the potential, as illustrated in Figure 8. In the metastable states for which $k=2\pi n/L$ and $n>1$ the circle is mapped $n$ times around the brim of the hat. As $n$ increases, the amplitude of the wave decreases as the kinetic term of the free energy acts as a tension in the line. When $|r|<k^{2}$, Eq. (19) no longer corresponds to a stationary solution of the equations of motion, neither stable, metastable nor unstable, as the tension has become too high. This solution is closely related to the topological defects this model admits [52], and the general analysis holds for small finite $\bar{\varphi}$, see video_2.mp4 in[55] for further illustration.

### v2

@@ IV.2 Continuous Symmetry
We can find an exact solution for the $\mathrm{SO}(2)$ model in the case of $\bar{\varphi}=0$, corresponding to a slice of the linear stability diagram of the last section where the entire plane, for $r<0$, is unstable [54, 51]. This solution has the form
@@
Decreasing further below $r=0$ yields shorter wavelength solutions. Minimizing the free energy density $\mathcal{F}=-\frac{3}{4}uA^{2}$ corresponds to maximizing the amplitude of the wave. The global minimum is the state with the longest wavelength $k=2\pi/L$ where $L$ is the system size, while solutions with higher wave numbers are metastable states, at least for $\alpha=0$. Assuming $|r|\gg k^{2}$, the amplitude of the wave is $A\approx\sqrt{-r/u}$, corresponding to mapping the circle in field space around the bottom of the potential, as illustrated in Figure 8. In the metastable states for which $k=2\pi n/L$ and $n>1$ the circle is mapped $n$ times around the brim of the hat. As $n$ increases, the amplitude of the wave decreases as the kinetic term of the free energy acts as a tension in the line. This solution is closely related to the topological defects this model admits [52], as far away from the core of a defect, the system forms a traveling sine-wave. The general analysis holds for small finite $\bar{\varphi}$, see video_2.mp4 in[56] for further illustration. When $|r|<k^{2}$, Eq. (19) no longer corresponds to a stationary solution of the equations of motion, neither stable, nor metastable, nor unstable, as the tension has become too high.
@@
Videos 2 and 3 are simulations of the NRCH model with an added noise term od the form $\sqrt{2D}\,\xi_{a}(\bm{x},t)$, where $\xi$ represents Gaussian white noise of unit strength with zero mean. The noise-strength in the simulations is set to $D=2\times 10^{-4}$ in the appropriate non-dimensionalized form. The value is chosen to show the robustness of the observed behavior to weak noise, while still illustrating the deviation away from the analytical solution of Figure 9.
IV.3 Relation to the Complex Ginzburg-Landau equation
We would like to highlight that a non-conserved version of our $\mathrm{SO}(2)$ model, which can be written as
[…]
has similarities to the Complex Ginzburg-Landau equation (CGLE) [29]. One may consider CGLE as the non-reciprocal extension of model A [65], and the $\mathrm{SO}(2)$ NRCH, Eq. (5), as the non-reciprocal extension of model B. The CGLE may include higher-order non reciprocal terms, which have been explored recently for the NRCH [54]. In the case of the CGLE, these term are necessary in order to obtain non-trivial behavior, as the liner term can be removed by a time-dependent field transformation, $\varphi_{a}\rightarrow[e^{-\alpha t\epsilon}]_{ab}\varphi_{b}$. At a formal level, one might, naively, extend this argument to the $\mathrm{SO(2)}$ NRCH as well, with the transformation $\varphi_{a}\rightarrow[e^{\alpha\nabla^{2}t\epsilon}]_{ab}\varphi_{b}$. Applying this transformation to the traveling-wave solution Eq. (19) gives a static-wave solution that solves the $\alpha=0$ equation of motion. This would hint at a hidden $\mathbb{Z}_{2}$ symmetry, so that the full symmetry is $\mathrm{O}(2)\simeq\mathrm{SO}(2)\rtimes\mathbb{Z}_{2}.$
We note, however, that the conservation law introduces intricacies into the structure of the equation that hampers the applicability of this argument, as the Laplacian operator in the exponential renders the transformation highly non-local. For example, the NRCH admits a broken-symmetry state in two dimensions [51], which is not the case case for model B. Furthermore, while Eq. (22) admits a Gaussian steady state, as one would expect since it can be mapped to an equilibrium theory, the same is not the case for Eq. (5). This indicates that the transformation is in general not well-defined.